Solve the formula for h: V=1/3 Bh
Thanks,
Pat
: Solve the formula for h: V=1/3 Bh
Thanks,
Pat

Answer by scott8148(2801)

multiplying by 3 __ 3V=Bh __ dividing by B __ 3V/B=h
Question 151232: Solve the formula for h: V=1/3 Bh
Thanks,
Pat
: Solve the formula for h: V=1/3 Bh
Thanks,
Pat

Answer by checkley77(3848)

V=1/3 Bh
V=BH/3 MULTIPLY BOTH SIDES BY 3.
BH=3V DIVIDE BOTH SIDES BY B.
H=3V/B ANSWER.